OpenStudy (anonymous):

Is there any easy way to figure out what the first step should be with Trig simplifying/proofs? Could I get a hint for the first step in cos(x)+sin(x) ----------- cos(2x)

OpenStudy (anonymous):

expand the bottom

OpenStudy (anonymous):

whenever u see double angles expand to become single angles

OpenStudy (anonymous):

So expand to \[\cos x + \sin x \div \cos^2x -\sin^2x\] ?

OpenStudy (anonymous):

The next step I did from that is take the reciprocal: cos^2(x) - sin^2(x) - --------------- cos(x) +sin(x) I intended to just divide out the exponent, but I'm 90% sure that's impossible because of the presence of the addition/subtraction.

OpenStudy (anonymous):

no, for starters that would get you going round in circles

OpenStudy (anonymous):

this one is kinda of easy and a little tricky

OpenStudy (anonymous):

you are pretty much finsihed with what you did before

OpenStudy (anonymous):

\[\frac{\cos(x) +\sin(x)}{\cos^2(x)-\sin^2(x)} \]

OpenStudy (anonymous):

follow from that

OpenStudy (anonymous):

So with that one, it's just the one step?

OpenStudy (anonymous):

no, thats not finished yet lol

OpenStudy (anonymous):

what if I said simplify \[\frac{u+3}{u^2 -9} \]

OpenStudy (anonymous):

if a very similar question

OpenStudy (anonymous):

(I'm retaking this math course, and this is exactly what I was struggling with the first time, I think I just don't know basic algebra which is why I can't do these very well D:) I'm not sure how to divide things out when there's additions and the numbers/variables don't have common multiples...

OpenStudy (anonymous):

its just an idenity that you must remember that (a^2 -b^2) = (a-b)(a+b)

OpenStudy (anonymous):

its not that hard to remember

OpenStudy (anonymous):

always looking for sqaure of numbers

OpenStudy (anonymous):

why you should know atleast up to your 12 times tables off you head